# Optimal Control on Positive Cones

### v1

@@ 1 Introduction
The theory of optimal control has been a cornerstone of control theory ever since the early work by Bellman [1]. Under standard assumptions, a discrete time optimal control problem with non-negative cost
Minimize $\sum_{t=0}^{\infty}g(x(t),u(t))$
subject to $x(t+1)=f(x(t),u(t)),\quad x(0)=x_{0}$
$u(t)\in U(x(t))$
is solvable if and only if the Bellman equation
$J^{*}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J^{*}(f(x,u))\right]$
has a non-negative solution $J^{*}$ which is finite for all $x$. The minimizing $u$ gives an optimal control policy. See for example [2, section 3.1].
Of particular importance are the rare situations when the Bellman equation has an explicit solution. The most well known case is probably linear dynamics and quadratic cost, introduced in the pioneering work of [3]. Another case of fundamental importance is when both $x$ and $u$ are restricted to finite sets. This setting leads to shortest path problems on graphs, with an extensive literature dating back to the work of [4]. Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$, i.e. so called positive systems, see [5, 6, 7].
The purpose of this note is to show that the previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
@@ 2 Main Result
Let $\mathcal{P}$ be a positive cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. Given $(s,r)\in\mathcal{P}^{*}$ and $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$. Moreover, assume the existence of $\phi:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that
@@
for all $(x,v)\in\mathcal{P}$, $(\lambda,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
@@
    The following optimal control problem has a finite value for every $(x_{0},u_{0})\in\mathcal{P}$:
    Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$
    subject to $x(t+1)=Ax(t)+Bu(t)$
    $(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
@@
    There exists $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$ (2)
@@
The minimization in ($i$) is a special case of the general formulation in the introduction, where $f(x,u)=Ax+Bu$ and $g(x,u)=s^{\top}x+r^{\top}u$. Define the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ recursively by
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}),$ $\lambda_{0}$ $:=0.$
If $J_{k}(x)=\lambda_{k}^{\top}x$, then
$\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right]$
@@
$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{k})^{\top}u\right]$
$=\left[s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})\right]^{\top}x$
$=\lambda_{k+1}^{\top}x.$
Hence, value iteration
$J_{k+1}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right],$ $J_{0}$ $=0$
gives a sequence $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$. In particular $(\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$. If ($i$) holds, then the sequence has an upper limit $J^{*}(x)=\lambda_{*}^{\top}x$ with $\lambda_{*}$ satisfying the condition in ($ii$). Conversely, if ($ii$) holds, then $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$, so the sequence $J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ has an upper limit and ($i$) holds too. This proves the equivalence of ($i$) and ($ii$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$). The optimal controller is $u(t)=\lambda(x(t)$, where
$\lambda(x)$ $=\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
$=\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right].$
It remains to prove equivalence betwen ($ii$) and ($iii$).
First, by definition of $\phi$, the optimization constraint in ($iii$) implies that
$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ (3)
for all $(x,u)\in\mathcal{P}$. It is a well known fact that solutions to the Bellman inequality are lower bounds on the optimal cost. Hence, if ($ii$) holds, the maximum for the convex program must be attained at the solution of eq. 2, so also ($iii$) holds. Conversely, suppose that ($iii$) holds with the maximizing argument $\lambda_{k}$. Define
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$
It follows from eq. 3 that
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)$
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+\lambda_{k}^{\top}Ax+\phi(r+B^{\top}\lambda_{k})^{\top}x$
for all $(x,u)\in\mathcal{P}$ and therefore
$(\lambda_{k+1}-\lambda_{k},0)$ $=(s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})-\lambda_{k},0)\in\mathcal{P}_{*}.$
It follows that
$\lambda_{k+1}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k+1}^{\top}(Ax+Bu)$
Hence also $\lambda_{k+1}$ satisfies the constraints of the optimization in ($iii$), so $\lambda_{k+1}=\lambda_{k}$. Otherwise $\lambda_{k}$ would not be maximizing. This means that $\lambda_{k}$ solves eq. 2 and also ($ii$) holds. The proof is complete. ∎
@@
1 supposes the existence of a function $\phi$ with the key property that it renders the optimization in eq. 1 linear in $x$. It is this feature that keeps the value iteration linear in $x$, which is instrumental in allowing the recasting of the optimal control problem as either a fixed point equation or convex program (conditions ($ii$) and ($iii$) in 1). It is natural then to ask when such a $\phi$ exists. The following result shows that when the cone $\mathcal{P}$ is proper, this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$. We will use this result in the next section when we connect 1 to some well studied optimal control problems.
@@
Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$, $\lambda^{*}\in\mathbb{R}^{n}$, and $\mu\in\mathbb{R}^{m}$. Suppose that there exists a $\lambda\in\mathbb{R}^{n}$ such that
$\left(\lambda,\mu\right)\in\operatorname{rel}\operatorname{int}\mathcal{P}^{*}.$ (4)
Then the following statements are equivalent:
@@
    $-\lambda^{*\top{}}x=\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u.$ (5)
@@
    $\left(\lambda^{*},\mu\right)\in\mathcal{P}^{*},\;\text{and for all}\;\left(\lambda,\mu\right)\in\mathcal{P}^{*},\;\left(\lambda-\lambda^{*},0\right)\in\mathcal{P}^{*}.$
@@
($i$$\implies$($ii$): With a view to generating a contradiction, suppose that under the hypothesis of ($i$),
$\left(\lambda^{*},\mu\right)\notin\mathcal{P}^{*}.$
There therefore exists an
$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda^{*}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
which implies that $-\lambda^{*\top{}}\bar{x}>\mu^{\top}\bar{u}$. However from eq. 5 we see that $-\lambda^{*\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction. This proves the first claim in ($ii$). Now, again with a view to generating a contradiction, suppose that
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\;\text{and}\;\left(\lambda-\lambda^{*},0\right)\notin\mathcal{P}^{*}.$
Therefore there exists an
$\left(\tilde{x},\tilde{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}\tilde{x}\\ \tilde{u}\end{bmatrix}<0,$
which implies that $\lambda^{*\top{}}\tilde{x}>\lambda^{\top}\tilde{x}$. The dual of eq. 5 reads
@@
The solution to eq. 6 is lower bounded by $-\lambda^{\top}\tilde{x}$. However by weak duality, the solution to eq. 6 is upper bounded by $-\lambda^{*\top{}}\tilde{x}$. Therefore
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\lambda^{*\top{}}\tilde{x}\leq{}\lambda^{\top}\tilde{x}.$
Again this a contradiction, proving the claim.
($ii$$\implies$($i$): Since $\mathcal{P}$ is a proper cone, $\mathcal{P}^{**}=\mathcal{P}$. Therefore for all $\left(x,u\right)\in\mathcal{P}$,
$\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\lambda^{*}\right)^{\top}x\leq{}0.$
@@
$-\lambda^{\top}x=-\left(\lambda-\lambda^{*}\right)^{\top}x-\lambda^{*\top{}}x\leq{}-\lambda^{*\top{}}x.$
It then follows that the solution to eq. 6 equals $-\lambda^{*\top{}}x$. Condition eq. 4 states that eq. 6 is strictly feasible. This implies that strong duality holds, which proves eq. 5 as required. ∎
@@ 3.1 The Linear Quadratic Regulator
$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succeq{}0.$ (7)
@@
having a minimum element (that is, an element $\Lambda^{*}\in\mathcal{C}_{\mu}$ such that $\Lambda^{*}\preceq{}\Lambda$ for all $\Lambda\in\mathcal{C}_{\mu}$). For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\Lambda^{*}=M_{1}M_{2}^{+}M_{1}^{\top}{}$. Therefore
@@
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Note that a degree of care has to be taken when reading this formula, since the requirements on the domain of $\phi$ have been left implicit. For example, these are guaranteed to hold under the common assumption that $R_{2}\succ{}0$, after which the above collapses to the usual algebraic Riccati equation. This can then be solved in the usual way whenever the pair $\left(F,G\right)$ is stabilizable. 1 ($iii$) can be similarly analyzed. In this case, the convex program reduces to the semi-definite program
@@
$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=$
$\qquad\qquad\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
That is, under eq. 8, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical. It then follows (with a little extra work) that the standard LQR problem corresponds to the special case of the matrix optimization when the initial condition $X_{0}$ is rank 1.

### v2

@@ Abstract
Notation
A cone $\mathcal{P}\subseteq{}\mathbb{R}^{n}$ is said to be proper if it is convex, closed, pointed and has a non-empty interior. The dual cone of a proper cone $\mathcal{P}$ is also proper, and given by
$\mathcal{P}^{*}=\left\{\lambda{}:\lambda{}^{\top}{}x\geq{}0\text{ for all }x\in\mathcal{P}\right\}.$
All further non-standard notation will be introduced as it is used.
@@ 1 Introduction
This paper is concerned with optimal control of the form
Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$ (1)
subject to $x(t+1)=Ax(t)+Bu(t)$
$(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
where $\mathcal{P}$ is a proper cone in $\mathbb{R}^{n}$. Assumptions are made to make sure that the objective function is non-negative and thus finite if and only if the Bellman equation
$J^{*}(x)$ $=\min_{(x(t),u(t))\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J^{*}(Ax+Bu)\right]$
has a non-negative solution $J^{*}$. The minimizing $u$ gives an optimal control policy. See for example [1, section 3.1].
The setting above turns out be one of the rare situations when the Bellman equation has an explicit solution. Otherwise, the most well known case from the literature is probably linear dynamics and quadratic cost, introduced in the pioneering work of [2]. Another case of fundamental importance is when both $x$ and $u$ are restricted to finite sets. This setting leads to shortest path problems on graphs, with an extensive literature dating back to the work of [3]. Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$ (the so called positive systems), see [4, 5]. A closely related case with bilinear dynamics is treated in [6].
The purpose of this note is to show that several previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
@@ 2 Main Result
Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. Given $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$. Let $(s,r)$ be an interior point of $\mathcal{P}^{*}$. Moreover, assume the existence of $\phi:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that
@@
for all $(x,v)\in\mathcal{P}$, $(\eta,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
@@
    The problem eq. 1 has a finite value for all $(x_{0},u_{0})\in\mathcal{P}$.
@@
    There exists a $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$
@@
Given $(\lambda_{0},0)\in\mathcal{P}^{*}$, consider the Bellman recursion
$J_{k+1}(x)$ $=\min_{x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J_{k}(Ax+Bu)\right],$
$J_{0}(x)$ $=\lambda_{0}^{\top}x$
and the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ generated by
$\lambda_{k+1}$ $=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$ (3)
We will prove by induction that this implies $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$: Assume for some $k\geq 0$ that $J_{k}(x)=\lambda_{k}^{\top}x$ and $(\lambda_{k},0)\in\mathcal{P}^{*}$. Let $(\eta_{k},\mu_{k}):=(s+A^{\top}\lambda_{k},r+B^{\top}\lambda_{k})$. Then
$\eta_{k}^{\top}x+\mu_{k}^{\top}u$ $=s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\geq 0$
for all $(x,u)\in\mathcal{P}$, so $(\eta_{k},\mu_{k})\in\mathcal{P}^{*}$. Hence
$J_{k+1}(x)$ $=\min_{x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J_{k}(Ax+Bu)\right]$
@@
$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[\mu_{k}^{\top}u\right]$
$=\left[s+A^{\top}\lambda_{k}+\phi(\mu_{k})\right]^{\top}x=\lambda_{k+1}^{\top}x$
and $\lambda_{k+1}^{\top}x\geq 0$, so $(\lambda_{k+1},0)\in\mathcal{P}^{*}$ and the induction assumption holds also for $k+1$. The assumption is trivial for $k=0$, so it follows by induction for $k=1,2,3,\ldots$
($ii$) $\Longleftrightarrow$ ($i$): If ($i$) holds, then the Bellman recursion from $J_{0}=0$ gives $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with upper limit $\lambda_{*}^{\top}x$ where $\lambda_{*}$ satisfies the condition in ($ii$). Conversely, if ($ii$), then $\lambda_{0}=0$ gives $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}^{*}$ for all $k$. This is trivial for $k=0$ and follows by induction for all other $k$ since
$\lambda_{k+1}^{\top}x$ $=\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\right]$
$\leq\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]=\lambda_{*}^{\top}x.$
Hence the $J_{k}$-sequence has an upper limit and ($i$) holds too. This proves the implication from ($ii$) to ($i$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$). Given $x$, the optimal control is $u$ is obtained by the minimization
$\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
or equivalently $\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right]$.
($ii$$\implies$($iii$): Consider $\lambda_{*}$ satisfying ($ii$) and let $\lambda$ be a feasible point for the optimization problem in ($iii$). In particular,
$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ for ($\text{for }(x,u)\in\mathcal{P}.$ (4)
The assumption that $(s,r)$ is an interior point of $\mathcal{P}^{*}$ gives existence of $\gamma,\delta>0$ such that
$\lambda_{*}^{\top}(Ax+Bu)$ $\leq\gamma(s^{\top}x+r^{\top}u)$
$\lambda^{\top}x$ $\leq\delta s^{\top}x\leq\delta\lambda_{*}^{\top}x$ for ($\text{for }(x,u)\in\mathcal{P}.$
Let $\{\lambda_{k}\}_{k=0}^{\infty}$ be the sequence generated by eq. 3 starting from $\lambda_{0}=\lambda$. Combination of eq. 3 and eq. 4 gives $(\lambda_{k}-\lambda,0)\in\mathcal{P}^{*}$ for all $k$. Moreover, [7, Proposition 2] implies that
00 0 $\leq\lambda_{k}^{\top}x\leq\left[1+\frac{\delta-1}{(1+\gamma^{-1})^{k}}\right]\lambda_{*}^{\top}x$ for all $\text{for all }k,$
so in the limit $(\lambda_{*}-\lambda,0)\in\mathcal{P}$ and $\lambda_{*}^{\top}x$ an upper bound for the value of the convex program in ($iii$).
($iii$$\implies$($ii$): Consider the sequence defined by eq. 3 initialized with $\lambda_{0}=0$. Then
$\lambda_{k}^{\top}x$ $\leq\lambda_{k+1}^{\top}x=s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu),$
so $\lambda_{k}$ is a feasible point for the optimization problem in ($iii$). Hence ($iii$) implies that the sequence $\lambda_{1}^{\top}x\leq\lambda_{2}^{\top}x\leq\ldots$ is bounded for every $x$. Therefore the limit $\lambda_{*}=\lim_{k\to\infty}\lambda_{k}$ exists and satisfies ($ii$). The proof is complete. ∎
@@
1 supposes the existence of a function $\phi$ with the key property that it renders the optimization in eq. 2 linear in $x$. It is this feature that keeps the value iteration linear in $x$, which is instrumental in allowing the recasting of the optimal control problem as either a fixed point equation or convex program (conditions ($ii$) and ($iii$) in 1). It is natural then to ask when such a $\phi$ exists. The following result shows that this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$. We will use this result in the next section when we connect 1 to some well studied optimal control problems.
@@
Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. For any $\mu\in\mathbb{R}^{m}$ such that the set
$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$ (5)
has a non-empty interior, the following are equivalent:
@@
    The function
    $x\mapsto\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u$ (6)
    with domain $\left\{x:\left(x,v\right)\in\mathcal{P}\text{ for some }v\right\}$ is a linear form (that is, the function above takes the form $f(x)=\phi^{\top}x$ for some $\phi\in\mathbb{R}^{n}$).
@@
    The set $\mathcal{C}_{\mu}$ has a minimum element (that is, there exists a $\bar{\lambda}\in\mathcal{C}_{\mu}$ such that $\lambda\in\mathcal{C}_{\mu}\implies\left(\lambda-\bar{\lambda},0\right)\in\mathcal{P}^{*}$).
Furthermore, if $\bar{\lambda}$ is the minimum element in ($ii$), then the linear form in ($i$) is $f(x)=-\bar{\lambda}^{\top}x$.
@@
($i$$\implies$($ii$): Under the hypothesis of ($i$), there exists a $\bar{\lambda}$ such that the function in eq. 6 takes the form $f\left(x\right)=-\bar{\lambda}^{\top}{}x$. We will first establish that $\left(\bar{\lambda},\mu\right)\in\mathcal{P}^{*}$. Assume the converse. There therefore exists an
$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\bar{\lambda}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
which implies that $-\bar{\lambda}^{\top}\bar{x}>\mu^{\top}\bar{u}$. However from eq. 6 we see that $-\bar{\lambda}^{\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction. We will now show that $\bar{\lambda}$ is the minimum element of $\mathcal{C}_{\mu}$. The optimization in eq. 6 is a cone program on standard form, with dual program (see for example [8, p.266])
@@
By weak duality this implies that
$\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u\geq{}\sup_{\left(\lambda,\mu\right)\in\mathcal{P}^{*}}-\lambda^{\top}x,$
and so for any $\left(x,u\right)\in\mathcal{P}$,
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\bar{\lambda}^{\top{}}x\leq{}\lambda^{\top}x.$ (8)
This implies that $\left(\lambda-\bar{\lambda},0\right)\in\mathcal{P}^{*}$, which proves the claim.
($ii$$\implies$($i$): For all $\left(x,u\right)\in\mathcal{P}$,
$\begin{bmatrix}\lambda-\bar{\lambda}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\bar{\lambda}\right)^{\top}x\leq{}0.$
@@
$-\lambda^{\top}x=-\left(\lambda-\bar{\lambda}\right)^{\top}x-\bar{\lambda}^{\top{}}x\leq{}-\bar{\lambda}^{\top{}}x.$
It then follows that the solution to eq. 7 equals $-\bar{\lambda}^{\top{}}x$. Condition eq. 5 ensures that eq. 7 is strictly feasible. Therefore strong duality holds, which proves eq. 6 as required. ∎
@@ 3.1 The Linear Quadratic Regulator
$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succ{}0.$ (9)
@@
having a minimum element. For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\overline{\Lambda}=M_{1}M_{2}^{+}M_{1}^{\top}{}$. Therefore
@@
$\Lambda_{*}$ $=S+F^{\top}\Lambda_{*}F-\left(F^{\top}\Lambda_{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda_{*}G\right)^{-1}\left(G^{\top}\Lambda_{*}F+R_{1}^{\top}{}\right).$
That is, the equation in 1($ii$) becomes the algebraic Riccati equation (the pseudo-inverses become inverses in light of eq. 9). This can then be solved in the usual way whenever the pair $\left(F,G\right)$ is stabilizable. 1($iii$) can be similarly analyzed. In this case, the convex program reduces to the semi-definite program
@@
$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
That is, under eq. 10, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical. It then follows (with a little extra work) that when the initial condition $X_{0}$ is rank 1, that the studied matrix optimization collapses to the standard LQR problem.